# Ultrafast All-Optical Measurement of Squeezed Vacuum in a Lithium Niobate Nanophotonic Circuit

### v2

@@ III Experimental Setup
Figure 4 shows the experimental setup. An 80 MHz mode-locked titanium-sapphire laser (MLL) sends 70-fs pulses to a half-wave plate (HWP) and polarizing beamsplitter (PBS) combination to control the power splitting between the measurement and squeezer beam paths. Pump light for the squeezer OPA is sent to a delay stage used to scan $\phi$ in equation 1. A shutter is also placed in this path to block the squeezer beam for shot-noise calibration. Two continuously variable neutral density filters (VND) are used in both paths to fine tune the input power. The paths are combined on a reflective objective (Obj) and focused onto the waveguide inputs for the OPAs. Figure 4A shows the layout of the nanophotonic circuit. In the squeezer OPA, the vacuum field around 1860 nm is squeezed by the pump. The SV and the remaining pump light enter an adiabatic coupler which passes at least 55% of the SV and 5% of the pump into the measurement OPA. Inside the measurement OPA, the SV is amplified by a strong pump pulse (40 pJ) to macroscopic photon levels detectable by a fast photodetector. A lensed fiber collects light from the output of the measurement OPA and directs it to a series of free-space filters used to remove remaining pump light, including the previously mentioned 1700-nm long pass and a 1950-nm short pass for limiting the measured mode number. From here, the filtered SV is collected into single-mode fiber and enters a 90/10 splitter where 90% is sent to a fast (10 GHz) photodetector and fast (80 GSPS 40 GHz) oscilloscope used to resolve individual pulses while the remaining 10% is sent to a slow (1MHz) detector and slow (100 MSPS 10 MHz) oscilloscope used to detect the amplification fringe and determine the measurement phase $\phi$.
@@ IV.1 Shot-noise Calibration
Initial characterization of the shot noise and squeezing levels is performed using a slow photodetector and an optical spectrum analyzer (OSA) which replaces the fast detector in Fig. 4. This data is used to calibrate the shot noise of the measurement OPA and measure the squeezing generated by the squeezer OPA (See [41] supplementary). During calibration, the mechanical shutter is initially left open. As the relative delay of the squeezer path is scanned, the measurement amplifier oscillates between measuring squeezed and anti-squeezed quadratures, resulting in the black curve in Fig. 5A. The remaining pump light collected from the output is monitored on a second OSA. Because the optical coupler between the measurement and squeezer OPAs couples some of the pump used to drive the squeezer, the measurement and squeezer pumps interfere inside the measurement OPA, causing the gain, and therefore shot-noise level, of the measurement OPA to depend on $\phi$. The shutter is then closed and the baseline shot noise is measured by the OSA resulting in the gray curve in Fig. 5A.
To determine the shot-noise levels corresponding to the minimum and maximum of the squeezing curve, the measurement OPA power is lowered using the VND until the measured pump power matches the minima of the pump interference. The shot-noise minimum is then measured on the slow photodetector and OSA resulting in the red curve in Fig. 5A. This same procedure is repeated for the shot-noise maximum as well to calibrate the anti-squeezing and obtain the blue curve. The pump OSA is set to measure around 970 nm to minimize any pump depletion effects. After shot-noise calibration, the squeezing is calculated as:
@@ IV.2 Slow Photodetector Measurements
Figure 5: A) Squeezing measurement taken on the OSA. The black curve is the amplification fringe produced from squeezed vacuum amplified by the measurement OPA. The grey curve is the baseline shot noise at the pump power used to measure the black curve. The red and blue curves are the corrected shot-noise levels to account for pump interference in the measurement OPA. B) Variance relative to vacuum vs squeezer pulse energy with a measurement pulse energy of 30 pJ. C) Variance relative to vacuum vs measurement pulse energy. Error bars are calculated from shot-noise variations during each measurement.
@@
Figure 5B shows the measured squeezing on the slow photodetector and OSA as a function of the input pulse energy for the squeezer waveguide with a measurement pulse energy of 30 pJ. For these measurements, the 1950-nm long-pass filter is removed to accentuate the impact of multiple modes (a Schmidt number of 2.8). Because the OSA uses a 2-nm bandpass filter, it can better suppress higher-order modes to isolate the squeezing in the fundamental mode [41, 53]. For the photodetector, the presence of higher-order modes with less squeezing reduces overall measured squeezing as the photodetector is not mode-sensitive and measures the squeezing as averaged over all modes. At low squeezer pulse energies, the measured squeezing is limited by the gain available in the squeezer OPA and grows as the pulse energy is increased. After 5 pJ, the observed squeezing begins to decrease with increasing pulse energy. Two phenomena contribute to this behavior. First, the finite phase noise of the pump laser causes a portion of the anti-squeezed quadrature to leak into the measurement of the squeezed quadrature. Raising the squeezer gain increases the contribution of the anti-squeezing to the measurement and reduces the measured squeezing. This increased gain produces more squeezing, but the noise reduction in the squeezed quadrature is asymptotically limited by the finite efficiency of the coupler while the noise of the anti-squeezed quadrature is not (see solid theory curves in Fig. 5B). Second, nonlinear behavior in the measurement amplifier increases with input power and further degrades measured squeezing.
@@
The solid lines plotted in Fig. 5B are calculated from the theoretical measured squeezing as a function of the coupling loss from the squeezer to measurement OPA. As SV passes through the coupler, some of the light is lost to propagation losses and fabrication imperfections in the coupler. This light is replaced with vacuum noise, causing lower squeezing and anti-squeezing values to be measured [63]. For a given measurement efficiency $\eta$, the measured squeezing can be modeled as
@@
Where $r$ is the on-chip squeezer gain in natural log units. Using this fit, we estimate an effective measurement efficiency ($\eta$) of 55% at 1860 nm. This fit includes squeezing lost to inefficiencies in our coupler as well as losses to nonlinear behavior in the measurement OPA. This fitting also allows us to calculate the expected anti-squeezing, however we can see that the measured data has a stark divergence from the expected trend. This is a result of the gain compression in the measurement OPA. Because the measurement OPA operates in the pump depleted regime, the anti-squeezed vacuum cannot be sufficiently amplified to reflect the true anti-squeezing entering the OPA. The same is true for the squeezing measurements in which the shot-noise level becomes compressed, leading to lower measured squeezing.
The effects of gain depletion can also be seen in Fig. 5C which plots the measured squeezing values vs the measurement pulse energy for a fixed squeezer energy of 5 pJ. As the measurement pulse energy is decreased, the measurement OPA experiences less gain depletion, leading to higher measured squeezing and anti-squeezing. In the undepleted pump regime, measured squeezing and anti-squeezing will follow Eq.10 and remain unchanged with small changes in pump energy. Hence, we know all measurements in Fig. 5B are taken in the pump depletion regime. This is also confirmed from Fig. 2C where pump depletion effects take hold around 20 pJ. Below 20 pJ, the shot noise no longer clears the electronic noise floor of the OSA, preventing squeezing measurements at lower measurement pump energies. This problem can be mitigated by improving output coupling efficiency and operating the measurement OPA at a lower pump power.
@@ IV.3 Pulse-to-pulse Measurements
Fast photodetector measurements resolve the pulse-to-pulse photon number and provide the statistical information necessary to measure $P(N,\phi)$ and recover the Wigner function. Before a fast measurement, a slow measurement is conducted with the OSAs to calibrate the shot noise. During each fast measurement, the shutter is initially left open, and the fast oscilloscope is triggered using the voltage applied to the piezo driving the squeezer’s delay stage to coincide with the middle of the $\phi$ scan. Data from the slow oscilloscope is used to determine $\phi$ and estimate the coupling stability based on measurement-to-measurement variations in the shot noise. After recording data, the squeezer’s shutter is closed and the measurement is repeated to find the calibrated shot-noise level.
@@
Figure 6 shows results gathered from fast photodetector measurements for a squeezer pulse energy of 5 pJ and a measurement pulse energy of 45 pJ. Sample measured histograms at 4 different phases are plotted in Fig. 6A, and each histogram corresponds to a single data point in Fig. 6B. For each measurement, we fit a 2-mode photon number distribution model to extract the variance of the fundamental mode as a function of $\phi$. Figure 6B shows these measurements for three cases: slow photodetector measurements, fast photodetector measurements, and the 2-mode fit. Slow measurement values are calculated from the slow oscilloscope trace using the method outlined in section IV.2. Because of the photon number offset resulting from pump depletion, the numerator and denominator of Eq.9 have a small but constant offset, resulting in the measured squeezing and anti-squeezing being underestimated. Fast measurement values are calculated using the variance of the photon numbers measured at the fast detector. This overcomes the offset problem and measures more squeezing and anti-squeezing, but these measurements are still contaminated with remaining contributions from higher-order modes. The 2-mode fit addresses this problem by isolating the variance as a function of $\phi$ for the fundamental mode, leading to higher measured squeezing and anti-squeezing. After performing this fit for all measurements, the distributions are re-sampled and fed into a maximum likelihood algorithm which calculates the Wigner functions for squeezing data (Fig. 6C). We extract a squeezing in the fundamental mode of 2.41 $\pm$ 0.34 dB and an anti-squeezing of 3.87 $\pm$ 0.61 dB. Figure 6E shows the recovered density matrix corresponding to the Wigner function plotted in Fig. 6C. We show the expected density matrix given the measured squeezing and anti-squeezing in Fig. 6D, and the absolute difference between the expected and recovered in Fig. 6F where we calculate a fidelity of 0.9998 $\pm$ 0.0001. While our fitting strategy limits us to states with known distributions, we can overcome this limitation with tighter bandpass filtering to better reject higher-order modes. Our losses between the chip and fiber (10 dB) and from our filter setup (6 dB) currently prevent us from tighter filtering as the remaining light is no longer strong enough to provide a sufficiently large signal-to-noise ratio at the fast photodetector.
@@ V Conclusion
We have demonstrated OPA-based Wigner tomography for squeezed vacuum in the ultrafast regime using dispersion-engineered TFLN. The low dispersion of our OPA design allows operation with ultra-short pulses to exceed the bandwidth of homodyne measurement and access a new regime of THz repetition-rate measurement and computation. With narrower bandpass filtering at the output of our OPA and modest improvements in detection efficiency, we can perform Wigner tomography on arbitrary quantum states encoded in a femtosecond optical pulses. We show that ultrafast nanophotonic OPAs can serve as quantum measurement devices, paving the way for ultrafast quantum information processing to be realized in a room-temperature chip-scale platform.

### v3

@@ III Experimental Setup
Figure 4 shows the experimental setup. An 80 MHz mode-locked titanium-sapphire laser (MLL) sends 70-fs pulses to a half-wave plate (HWP) and polarizing beamsplitter (PBS) combination to control the power splitting between the measurement and squeezer beam paths. Pump light for the squeezer OPA is sent to a delay stage used to scan $\phi$ in equation 1. A shutter is also placed in this path to block the squeezer beam for shot-noise calibration. Two continuously-variable neutral density filters (VND) are used in both paths to fine-tune the input power. The paths are combined on a reflective objective (Obj) and focused onto the waveguide inputs for the OPAs. Figure 4A shows the layout of the nanophotonic circuit. In the squeezer OPA, the vacuum field around 1860 nm is squeezed by the pump to generate squeezed vacuum (SV). The SV and the remaining pump light enter an adiabatic coupler which passes at least 55% of the SV and 5% of the pump into the measurement OPA. Inside the measurement OPA, the SV is amplified by a strong pump pulse (40 pJ) to macroscopic photon levels detectable by a fast photodetector. A lensed fiber collects light from the output of the measurement OPA and directs it to a series of free-space filters used to remove remaining pump light, including the previously mentioned 1700-nm long pass and a 1950-nm short pass for limiting the measured mode number. From here, the filtered SV is collected into single-mode fiber and enters a 90/10 splitter where 90% is sent to a fast (10 GHz) photodetector and fast (80 GSPS 40 GHz) oscilloscope used to resolve individual pulses while the remaining 10% is sent to a slow (1 MHz) detector and slow (100 MSPS 10 MHz) oscilloscope used to detect the amplification fringe and determine the measurement phase $\phi$.
@@ IV.1 Shot-noise Calibration
To calibrate our squeezing and anti-squeezing measurements, we measure the $\phi$-dependent shot noise using the slow (1 MHz) photodetector and an optical spectrum analyzer (OSA) which replaces the fast detector in Fig. 4. The coupler in Fig. 4A couples some of the 930-nm pump light used to drive the squeezer into the measurement OPA, causing it to interfere with the pump light sent into the upper port of Fig. 4A. Because of this interference, the gain inside the measurement OPA, and hence the shot noise, is dependent on $\phi$.
Data from the shot-noise calibration procedure is show in figure 5A. During calibration, the mechanical shutter is initially left open. As the relative delay of the squeezer path is scanned, the measurement amplifier oscillates between measuring squeezed (with destructive pump interference) and anti-squeezed (with constructive pump interference) quadratures, resulting in the black curve in Fig. 5A. The remaining pump light collected from the output is monitored on a second OSA to determine the strength of the pump interference. The shutter is then closed and the baseline shot noise is measured by the OSA resulting in the gray curve in Fig. 5A. To find the shot-noise level corresponding to destructive pump interference (i.e. the shot-noise minimum), the intensity of the measurement OPA pump is lowered until the measured power at the pump OSA matches the minima of the previously measured pump interference. The shot-noise minimum is then measured on the slow photodetector and OSA resulting in the red curve in Fig. 5A. This same procedure is repeated for the shot-noise maximum to calibrate the anti-squeezing and obtain the blue curve. The pump OSA is set to measure around 970 nm to minimize any pump depletion effects. After shot-noise calibration, the squeezing is calculated as:
@@ IV.2 Slow Photodetector Measurements
Figure 5: A) Squeezing measurement taken on the OSA. The black curve is the amplification fringe produced from squeezed vacuum amplified in the measurement OPA. The grey curve is the baseline shot noise at the pump power used to measure the black curve. The red and blue curves are the corrected shot-noise levels to account for pump interference in the measurement OPA. B) Variance relative to vacuum vs squeezer pulse energy with a measurement pulse energy of 30 pJ. C) Variance relative to vacuum vs measurement pulse energy. Error bars are calculated from shot-noise variations during each measurement.
@@
Figure 5B shows the measured squeezing on the slow photodetector and OSA as a function of the input pulse energy for the squeezer waveguide with a measurement pulse energy of 30 pJ. For these measurements, the 1950-nm long-pass filter is removed to accentuate the impact of multiple modes (a Schmidt number of 2.8). Because the OSA uses a 2-nm bandpass filter, it can better suppress higher-order modes to isolate the squeezing in the fundamental mode [18, 33]. For the photodetector, the presence of higher-order modes with less squeezing reduces overall measured squeezing as the photodetector is not mode-sensitive and measures the squeezing as averaged over all modes.
At low squeezer pulse energies, the measured squeezing is limited by the gain available in the squeezer OPA and grows as the pulse energy is increased. After 5 pJ, the observed squeezing begins to decrease with increasing pulse energy. Two phenomena contribute to this behavior. First, the finite phase noise of the pump laser causes a portion of the anti-squeezed quadrature to leak into the measurement of the squeezed quadrature. Raising the squeezer gain increases the contribution of the anti-squeezing to the measurement and reduces the measured squeezing. This increased gain produces more squeezing, but the noise reduction in the squeezed quadrature is asymptotically limited by the finite efficiency of the coupler while the noise of the anti-squeezed quadrature is not (see solid theory curves in Fig. 5B). Second, gain saturation in the measurement amplifier increases with input power and further degrades measured squeezing.
The effects of gain saturation on measured squeezing can be more easily seen in Fig. 5C which plots the squeezing values vs the measurement pulse energy for a fixed squeezer energy of 5 pJ. As the measurement pulse energy is decreased, the behavior of the measurement OPA becomes more linear with less gain compression, leading to higher measured squeezing and anti-squeezing. In the undepleted pump regime, measured squeezing and anti-squeezing will follow Eq.10 and remain unchanged with small changes in pump energy. Hence, we know all measurements in Fig. 5B are taken in the pump depletion regime as the measured anti-squeezing deviates from the theoretical model calculated using the efficiency of the coupler. This is also confirmed from Fig. 3C where pump depletion effects take hold around 20 pJ. Below 20 pJ, the shot noise no longer clears the electronic noise floor of the OSA, preventing squeezing measurements at lower measurement pump energies. For this reason, data points in Fig. 5B are taken at a pump energy of 30pJ to improve SNR and shrink vertical error bars while sacrificing some measurable squeezing (3 dB at 30 pJ pump compared to 4.1 dB at 20 pJ pump in Fig. 5B and C respectively).
@@
The solid lines plotted in Fig. 5B are calculated from the theoretical measured squeezing as a function of the coupling loss from the squeezer to measurement OPA. As SV passes through the coupler, some of the light is lost as a result of fabrication imperfections in the coupler. This light is replaced with vacuum noise, causing lower squeezing and anti-squeezing values to be measured [39]. For a measurement efficiency $\eta$, the measured squeezing can be modeled as:
@@
where $r$ is the on-chip squeezer gain in natural log units. Using this fit, we estimate an effective $\eta$ of 55% at 1860 nm. This fit includes squeezing lost to inefficiencies in our coupler as well as losses to nonlinear behavior in the measurement OPA. We expect $\eta$ to be dominated by the coupler losses given the characterization measurements taken previously of this device [18]. The vacuum noise injected from losses contaminates the SV, resulting in a state with reduced squeezing and anti-squeezing measured by the measurement OPA.
Equation 10 also allows us to calculate the expected anti-squeezing, however, the measured data has a stark divergence from the expected trend. This is a result of gain-compression in the measurement OPA. Because the measurement OPA operates in the pump-depleted regime, the anti-squeezed vacuum cannot be sufficiently amplified to reflect the true anti-squeezing entering the OPA. The same is true for the squeezing measurements as the shot-noise level becomes compressed, leading to lower measured squeezing.
@@ IV.3 Pulse-to-pulse Measurements
Fast photodetector measurements resolve the pulse-to-pulse intensity and provide the statistical information necessary to resolve $P(N,\phi)$ and recover the Wigner function. During a fast measurement, the shutter is initially left open, and the fast oscilloscope is triggered using the voltage applied to the piezo driving the squeezer’s delay stage to coincide with the middle of the $\phi$ scan. Data from the slow oscilloscope is used to determine $\phi$ and estimate the coupling stability based on measurement-to-measurement variations in the shot noise. After recording data, the shutter is closed to block the squeezer and the measurement is repeated to find the calibrated shot-noise level.
@@
Figure 6 shows results gathered from fast photodetector measurements for a squeezer pulse energy of 5 pJ and a measurement pulse energy of 45 pJ. This measurement pulse energy was chosen to satisfy SNR requirements at the fast photodetector given the loss between the chip and detector. Sample measured histograms at 4 different phases are plotted in Fig. 6A. Each histogram corresponds to a single data point in Fig. 6B. For each measurement, we fit a 2-mode photon number distribution model to extract the variance of the fundamental mode as a function of $\phi$. Figure 6B shows these measurements for three cases: slow photodetector measurements, fast photodetector measurements, and the 2-mode fit. Slow measurement values are calculated from the slow oscilloscope trace using the method outlined in section IV.2. Because of the photon number offset resulting from pump depletion, the numerator and denominator of Eq.9 have a small but constant offset, resulting in the measured squeezing and anti-squeezing being underestimated. Fast measurement values are calculated using the variance of the photon numbers measured at the fast detector. This overcomes the offset problem and measures more squeezing and anti-squeezing, but these measurements are still contaminated with remaining contributions from higher-order modes. The 2-mode fit addresses this problem by isolating the variance as a function of $\phi$ for the fundamental mode, leading to higher measured squeezing and anti-squeezing. After performing this fit for all measurements, the distributions are re-sampled and fed into a maximum likelihood algorithm which calculates the Wigner function (Fig. 6C). We extract a squeezing in the fundamental mode of 2.41 $\pm$ 0.34 dB and an anti-squeezing of 3.87 $\pm$ 0.61 dB. Figure 6E shows the recovered density matrix corresponding to the Wigner function plotted in Fig. 6C. We show the expected density matrix given the measured squeezing and anti-squeezing in Fig. 6D, and the absolute difference between the expected and recovered in Fig. 6F where we calculate a fidelity of 0.9998 $\pm$ 0.0001. This fidelity serves as a sanity-check to ensure our data analysis is not introducing distortions into the measurement. While our fitting strategy limits us to states with known distributions, we can overcome this limitation with tighter bandpass filtering to better reject higher-order modes. Our losses between the chip and fiber (10 dB) and from our filter setup (6 dB) currently prevent us from tighter filtering as the remaining light is no longer intense enough to provide a sufficient signal-to-noise ratio at the fast photodetector.
@@ V Conclusion
We have demonstrated OPA-based Wigner tomography for squeezed vacuum in the ultrafast regime using dispersion-engineered TFLN. This is, to the extend of the knowledge of the authors, the first demonstration of all-optical quantum state tomography on-chip. The low dispersion of our OPA design allows operation with ultra-short pulses to exceed the bandwidth of traditional measurement techniques and access a new regime of THz repetition-rate measurement and computation. Achieving THz clock rates is a fundamental advancement for time-multiplexed systems such as cluster states and measurement-based quantum computation as higher clock rates translates to faster computation and allows for larger quantum circuits to be realized.
Our THz claim is based on dispersion calculations derived from the waveguide geometry measured using atomic force microscopy. We also note that similar dispersion-engineered OPAs with faster clock rates than the one presented here have been experimentally characterized [40]. Speed limitations imposed by the photodetector bandwidth are present in our experiment, but can be overcome with fast demultimplexing schemes which we elaborate in the supplementary.
Because we only measure intensity, our current experimental implementation is limited to tomography of states with zero mean field. However, OPA-based state tomography can be generalized to any arbitrary state by implementing either homodyne detection or a displacement scheme at the output, both of which are mature technologies [41, 12, 42, 43, 44]. As discussed in earlier sections, experimental limitations stemming from high losses between the chip and fast photodetector introduce measurement distortion from pump depletion and multimode effects. Both of these problems can be solved with narrower bandpass filtering and modest improvements in detection efficiency to allow for the measurement OPA to operate in the undepleted pump regime.
To conclude, we show that dispersion-engineered nanophotonic OPAs can serve as quantum measurement devices, paving the way for ultrafast quantum information processing to be realized in a room-temperature chip-scale platform.